### v3

# Extremal problems and the combinatorics of sumsets

Abstract.
This is a survey of old and new problems and results in additive number theory.
Key words and phrases:
Minimal basis, maximal nonbasis, oscillations of bases, Sidon set, Sidon system, $\varphi$-Sidon system, $B_{h}[g]$-set, linear form, representation function.
2010 Mathematics Subject Classification:
11-02,11B05,11B13, 11B34, 11B75, 11P70, 05A18
Supported in part by grant # 66197-00 54 from the PSC-CUNY Research Award Program.
1. Sumsets and asymptotic bases
Here are the best known theorems in additive number theory. Every nonnegative integer is the sum of four squares (Lagrange). Every nonnegative integer is the sum of nine nonnegative cubes (Wieferich), every sufficiently large integer is the sum of seven nonnegative cubes (Linnik), and almost every nonnegative integer is the sum of four nonnegative cubes (Davenport). For all $k\geq 2$ there is a number $g(k)$ such that every nonnegative integer is the sum of $g(k)$ $k$th powers (Hilbert). Every sufficiently large odd integer is the sum of three primes (Vinogradov). These results are statements about sums of a fixed number of elements chosen from an “interesting” set of nonnegative integers. For a survey of these results, see Nathanson [51].
The classical theorems motivate the following fundamental definitions. Let $A,A_{1},\ldots,A_{h}$ be finite or infinite subsets of the set $\mathbf{N}_{0}=\{0,1,2,3,\ldots\}$ of nonnegative integers. Let $\mathbf{N}=\{1,2,3,\ldots\}$ be the set of positive integers. We define the sumset
$A_{1}+\cdots+A_{h}=\{a_{1}+\cdots+a_{h}:a_{i}\in A_{i}\text{ for all }i=1,2,\ldots,h\}$
and the $h$-fold sumset
$hA=\underbrace{A+\cdots+A}_{\text{$h$ summands}}=\{a_{1}+\cdots+a_{h}:a_{i}\in A\text{ for all }i=1,2,\ldots,h\}.$
We also consider sums of subsets of the set $\mathbf{Z}^{n}$ of lattice points, of the the set $\mathbf{Z}/m\mathbf{Z}$ congruence classes modulo $m$, and, more generally, of any additive abelian semigroup.
The set $A$ of nonnegative integers is a basis of order $h$ if every nonnegative integer can be represented as the sum of $h$ not necessarily distinct elements of $A$, that is, if $hA=\mathbf{N}_{0}$. The set $A$ is an asymptotic basis of order $h$ if $hA$ contains all sufficiently large integers, or, equivalently, if $hA\sim\mathbf{N}_{0}$. (We write $X\sim Y$ if the symmetric difference $(X\setminus Y)\cup(Y\setminus X)$ is finite.) Thus, Lagrange’s theorem states that the set $A^{(2)}=\{k^{2}:k\in\mathbf{N}_{0}\}$ is a basis of order 4, Wieferich’s theorem states that the set $A^{(3)}=\{k^{3}:k\in\mathbf{N}_{0}\}$ is a basis of order 9, and Linnik’s theorem states that $A^{(3)}$ is an asymptotic basis of order 7.
Let $A$ be a set of nonnegative integers. The $h$-fold representation function $r_{h,A}(n)$ counts the number of representations of $n$ as a sum of $h$ elements of $A$, that is, the number of $h$-tuples
$(a_{1},\ldots,a_{h})\in A^{h}$
such that
$n=a_{1}+a_{2}+\cdots+a_{h}$
and
$a_{1}\leq a_{2}\leq\cdots\leq a_{h}.$
A central unsolved problem in additive number theory is the following conjecture of Erdös and Turän [23].
Problem 1.
Let $A$ be a set of nonnegative integers. If $A$ is an asymptotic basis of order $h$, then the representation function $r_{h,A}(n)$ is unbounded. Equivalently,
$\liminf_{n\rightarrow\infty}r_{h,A}(n)>0$
implies
$\limsup_{n\rightarrow\infty}r_{h,A}(n)=\infty.$
One approach to the Erdős-Turán conjecture is to look for extremal asymptotic bases, that is, bases that are, in some sense, “small” and might have small representation functions. We consider two kinds of extremal bases: Thin bases and minimal bases. This paper discusses old and new results and old and new problems related to these extremal constructions in combinatorial and additive number theory.
The Erdös-Turän conjecture was originally stated for bases of order 2, but extends naturally to bases of all orders $h\geq 2$. Many theorems in additive number theory about 2-fold sumsets do not extend or have not (yet) been extended to $h$-fold sumsets for $h\geq 3$. One difficulty that often arises in combinatorial and probabilistic arguments is the lack of independence of representations. Two representations
$n=a_{1}+\cdots+a_{h}\operatorname{\qquad\text{and}\qquad}n=a^{\prime}_{1}+\cdots+a^{\prime}_{h}$
are independent if
$\{a_{1},\ldots,a_{h}\}\cap\{a^{\prime}_{1},\ldots,a^{\prime}_{h}\}=\emptyset.$
All representations with $h=2$ summands are independent: If $n=a_{1}+a_{2}=a^{\prime}_{1}+a^{\prime}_{2}$, then either $\{a_{1},a_{2}\}=\{a^{\prime}_{1},a^{\prime}_{2}\}$ or $\{a_{1},a_{2}\}\cap\{a^{\prime}_{1},a^{\prime}_{2}\}=\emptyset$. This is not true for representations as sums of more than two elements; for example, $8=1+2+5=1+3+4$. This phenomenon, known for a long time, has recently been rebranded as “additive energy.”
In the study of $h$-fold sumsets of a set $A$, we define the solution set $S_{h,A}(n)$:
$S_{h,A}(n)=\{a_{1}\in A:\text{there exist $a_{2},\ldots,a_{h}\in A$ such that $n=a_{1}+a_{2}+\cdots+a_{h}$}\}.$
Let $\Omega$ be the space of all infinite sets of positive integers, and let $p_{n}\in[0,1]$ for all $n\in\mathbf{N}$. Let
$E_{n}=\{A\in\Omega:n\in A\}.$
Erdős and Rényi [22] constructed a probability measure $\mu$ on $\Omega$ such that $\mu(E_{n})=p_{n}$ and the events $E_{1},E_{2},E_{2},\ldots$ are independent. We have the following result about solution sets.
Theorem 1 (Erdős-Nathanson [20]).
For $\alpha,\beta,\gamma\in\mathbf{R}$ with $\alpha>0$ and $1/3<\gamma\leq 1/2$, let
$p_{n}\leq\frac{\alpha\log^{\beta}(n+1)}{n^{\gamma}}$
for all $n\in\mathbf{N}$. For almost all $A\in\Omega$ and for all but finitely many pairs $(m,n)$ with $m\neq n$,
$|S_{2,A}(m)\cap S_{2,A}(n)|\leq\frac{2}{3\gamma-1}.$
Choosing $\gamma=1/2$, we obtain sets $A$ such that $|S_{2,A}(m)\cap S_{2,A}(n)|\leq 4$ for all but finitely many pairs $(m,n)$ of integers $m\neq n$.
Problem 2.
Extend Theorem 1 to solution sets $S_{h,A}(n)$ for $h\geq 3$.
2. Thin bases
Our first measure of smallness is “slow growth.” Let $A$ be a set of nonnegative integers. The counting function $A(x)$ counts the number of elements of $A$ that are no greater than $x$, that is,
$A(x)=\sum_{\begin{subarray}{c}a\in A\\ a\leq x\end{subarray}}1.$
If $A$ is an asymptotic basis of order $h$, then $r_{h,A}(n)\geq 1$ for all $n\geq n_{0}$. If $n_{0}\leq n\leq x$ and $n=a_{1}+\cdots+a_{h}\in hA$, then $a_{i}\leq x$ for all $i\in\{1,\ldots,h\}$. The number of $h$-tuples $(a_{1},\ldots,a_{h})\in A^{h}$ with $a_{1}\leq\cdots\leq a_{h}\leq x$ is $\binom{A(x)+h-1}{A(x)-1}=\binom{A(x)+h-1}{h}$. It follows that
$x-n_{0}\leq\binom{A(x)+h-1}{h}$
and so
$A(x)\gg x^{1/h}.$
An asymptotic basis $A$ of order $h$ is thin if
$A(x)\ll x^{1/h}.$
There are simple constructions of thin bases due to Raikov [70], Jia [35, 36, 37], and Nathanson [47, 49, 54]. Shatrovskii [6, 54] constructed beautiful examples of thin bases and, for every $h\geq 2$, Cassels [5, 53] constructed a thin asymptotic basis $A=\{a_{k}:k=1,2,\ldots\}$ such that $a_{k}=ck^{h}+O\left(k^{h-1}\right)$ for some $c>0$.
An asymptotic basis of order $h$ does not necessarily contain a subset that is a thin basis of order $h$, but it is natural to ask if the classical bases in additive number theory contain thin bases.
The set $A^{(2)}=\{k^{2}:k\in\mathbf{N}_{0}\}$ is a basis of order 4 with counting function $A(x)\sim x^{1/2}$. A basis of order 4 must have counting function $\gg x^{1/4}$ and so the set of all squares is “thick”. It is an unsolved problem to determine if there is a set of squares that is a thin basis of order 4. Probability arguments do show that there exist sets $W$ of nonnegative integers of density 0 such that the set of squares $\{w^{2}:w\in W\}$ is a basis of order 4. No explicit example is known. For work on this problem, see Choi-Erdős-Nathanson [10], Erdős-Nathanson [19], Wirsing [79], and Zöllner [80, 81].
Problem 3.
Construct a set $W$ of nonnegative integers of density 0 such that the set $\{w^{2}:w\in W\}$ is a basis of order 4.
There is the analogous problem for sums of $k$th powers (Waring’s problem). Hardy and Littlewood proved that for every $k\geq 2$ there is an integer $s_{0}(k)$ such that, if $s\geq s_{0}(k)$, then a positive integer has $cn^{(s/k)-1}$ representations as a sum of $k$th powers (Davenport [11], Vaughan [77]). This is used to prove the following.
Theorem 2 (Nathanson [48]).
For any $s>s_{0}(k)$ and $0<\varepsilon<1/s$, there is a probability measure on the set $\Omega$ of all infinite sets of positive integers such that, with probability 1, a random set $W\in\Omega$ has the following properties:
  1. (i)
    $W(x)\sim cx^{1-(1/s)+\varepsilon}$ for some $c>0$,
  2. (ii)
    every sufficiently large integer $n$ can be repsented in the form
    $n=w_{1}^{k}+2_{2}^{k}+\cdots+w_{s}^{k}$
    with $w_{1},w_{2},\ldots,w_{s}\in W$.
Van Vu [78] has improved this result.
Problem 4.
Construct a set $W_{k}$ of nonnegative integers of density 0 such that the set $\{w^{k}:w\in W_{k}\}$ is a basis of order h for some $h$.
3. Minimal asymptotic bases
A second measure of the smallness of an asymptotic basis is “minimality”. An asymptotic basis $A$ of order $h$ is minimal if no proper subset of $A$ is an asymptotic basis of order $h$. This means that, for every $a\in A$, the set $A\setminus\{a\}$ is not an asymptotic basis of order $h$, and so $h\left(A\setminus\{a\}\right)$ is a co-infinite subset of $\mathbf{N}_{0}$. Thus, for every $a\in A$, there are infinitely many nonnegative integers $n$, all of whose representations as sums of $h$ elements of $A$ must include the integer $a$.
We ask the following questions: Do minimal asymptotic bases exist, and, if so, does every asymptotic basis contain a minimal asymptotic nonbasis? In fact, minimal asymptotic bases of order $h$ do exist for all $h\geq 2$, but there are few explicit constructions. Nathanson [47, 49] used a 2-adic construction to produce the first examples of minimal asymptotic bases. This method was extended to $g$-adically defined sets by Chen [8], Chen and Chen [7], Chen and Tang [9], Jia [37], Jia and Nathanson [38], Lee [41], Li and Li [42], Ling and Tang [43, 44], Sun [73, 74], and Sun and Tao [75]. In these examples, one has $r_{h,A}(n)=1$ for infinitely many $n$, but also $\limsup_{n\rightarrow\infty}r_{h,A}(n)=\infty$, so the known examples do not provide counterexamples to the Erdős-Turän conjecture.
Nathanson [63] recently constructed a new class of minimal asymptotic bases of order $h$ for all $h\geq 2$. A $\mathcal{G}$-adic sequence is a strictly increasing sequence of positive integers $\mathcal{G}=(g_{j})_{j=0}^{\infty}$ such that $g_{0}=1$ and $g_{j-1}$ divides $g_{j}$ for all $j\geq 1$. Let $W$ be a nonempty set of nonnegative integers, and let $\mathcal{F}^{*}(W)$ be the set of all nonempty finite subsets of $W$. Let $\mathcal{G}=(g_{j})_{j=0}^{\infty}$ be a $\mathcal{G}$-adic sequence with $d_{j+1}=g_{j+1}/g_{j}$. We define the set of positive integers
$A_{\mathcal{G}}(W)=\left\{\sum_{j\in F}x_{j}g_{j}:F\in\mathcal{F}^{*}(W)\text{ and }x_{j}\in[1,d_{j+1}-1]\right\}.$
Theorem 3 (Nathanson [63]).
Let $h\geq 2$ and let $t$ be an integer such that
$t\geq 1+\frac{\log h}{\log 2}.$
Let $\mathcal{W}=(W_{i})_{i=0}^{h-1}$ be a partition of $\mathbf{N}_{0}$ such that, for all $i\in[0,h-1]$, the set $W_{i}$ contains infinitely many intervals of $t$ consecutive integers. Let $\mathcal{G}$ be a $\mathcal{G}$-adic sequence. The set
$A_{\mathcal{G}}(\mathcal{W})=\bigcup_{i=0}^{h-1}A_{\mathcal{G}}(W_{i})$
is a minimal asymptotic basis of order $h$.
Nathanson [47] proved that not every asymptotic basis contains a minimal asymptotic basis. There is the following remarkable example.
Theorem 4 (Erdős-Nathanson [15]).
There exists an asymptotic basis $A$ of order 2 such that, if $S$ is a subset of $A$, then $A\setminus S$ is an asymptotic basis of order 2 if and only if the set $S$ is finite.
Problem 5.
Let $h\geq 3$. Construct an asymptotic basis $A$ of order $h$ such that, if $S$ is a subset of $A$, then $A\setminus S$ is an asymptotic basis of order $h$ if and only if the set $S$ is finite.
It is a difficult problem to determine if an asymptotic basis contains a minimal asymptotic basis. Here is a result that uses the representation function.
Theorem 5 (Erdős-Nathanson [18]).
Let $A$ be an asymptotic basis of order $2$. If
(1) $r_{2,A}(n)>\frac{1}{\log(4/3)}\log n$
for all sufficiently large $n$, then $A$ contains a minimal asymptotic basis of order 2.
Problem 6.
The dependence on the particular constant $1/\log(4/3)$ in Theorem 5 is not understood. It is open problem to determine if the condition
$r_{2,A}(n)>c\log n$
for some $c>0$ and all sufficiently large $n$ implies that $A$ contains a minimal asymptotic basis of order 2.
Problem 7.
Determine if the condition
$\lim_{n\rightarrow\infty}r_{2,A}(n)=\infty$
implies that $A$ contains a minimal asymptotic basis of order 2.
Problem 8.
Let $h\geq 3$. Does there exists a constant $c_{h}>0$ such that the condition
$r_{h,A}(n)>c_{h}\log n$
for all sufficiently large $n$ implies that $A$ contains a minimal asymptotic basis of order $h$?
Inequality (1) appears mysteriously in the following result about partitioning an asymptotic basis into a disjoint union of asymptotic bases.
Theorem 6 (Erdős-Nathanson [21]).
Let $A$ be an asymptotic basis of order 2 whose representation function satisfies (1). There are subsets $A_{1}$ and $A_{2}$ of $A$ such that $A=A_{1}\cup A_{2}$ and $A_{1}\cap A_{2}=\emptyset$ and both $A_{1}$ and $A_{2}$ are asymptotic bases of order $h$.
Again, the dependence on the constant $1/\log(4/3)$ is not understood.
4. Maximal asymptotic nonbases
Minimal asymptotic bases are extremal objects in additive number theory.
Nathanson [47] introduced the dual notion. A set of nonnegative integers is an asymptotic nonbasis of order $h$ if it is not an asymptotic basis of order $h$. An asymptotic nonbasis $A$ of order $h$ is maximal if $A\cup\{a\}$ is an asymptotic basis of order $h$ for every nonnegative integer $a\notin A$. We ask the following questions: Do maximal asymptotic nonbases exist, and, if they do, is every asymptotic nonbasis contained in a maximal asymptotic nonbasis?
There are trivial examples of maximal asymptotic nonbases. For example, the set of even integers is a maximal asymptotic nonbasis of order 2. Other simple examples can be constructed that are unions of congruence classes. There are also nontrivial examples. Erdős and Nathanson [15] constructed a class of maximal asymptotic nonbases of order 2 that contain arbitrarily long gaps and so are not unions of congruence classes.
The first example of an asymptotic nonbasis that is not contained in a maximal asymptotic nonbasis is due to Hennefeld [34], who constructed a set $A$ with the property that a superset $A^{\prime}$ of $A$ is a nonbasis if and only if $A^{\prime}$ is a co-infinite subset of $\mathbf{N}_{0}$. There exists no maximal co-infinite subset of an infinite set, and so $A$ is not contained in a maximal nonbasis.
Here is a recent construction of a new class of asymptotic nonbases that are not contained in maximal asymptotic nonbases. The set $Y$ of integers has infinite gaps if, for all $C>0$, there exist only finitely many pairs of integers $y,y^{\prime}\in Y$ with $y\neq y^{\prime}$ and $|y-y^{\prime}|\leq C$.
Theorem 7 (Nathanson [56]).
Let $h\geq 2$. Let $s$ and $t$ be integers such that $\gcd(h,s-t)=1$.
  1. (a)
    Let $Y$ be an infinite set of integers with infinite gaps, and let $X=\mathbf{Z}\setminus Y.$ The set
    $A_{X}=\{s\}\cup\{hx+t:x\in X\}$
    is an asymptotic nonbasis of order $h$ for $\mathbf{Z}$ that is not a subset of a maximal asymptotic nonbasis of order $h$ for $\mathbf{Z}$.
  2. (b)
    Let $Y_{0}$ be an infinite set of nonnegative integers with infinite gaps, and let $X_{0}=\mathbf{N}_{0}\setminus Y_{0}$. The set
    $A_{X_{0}}=\{s\}\cup\{hx+t:x\in X_{0}\}$
    is an asymptotic nonbasis of order $h$ for $\mathbf{N}_{0}$ that is not a subset of a maximal asymptotic nonbasis of order $h$ for $\mathbf{N}_{0}$}.
5. Oscillations
We can extend the notion of minimal basis and maximal nonbasis by considering perturbations of a set $A$ of nonnegative integers by a finite set $F$ of integers with more than one element. Let $A$ be an asymptotic basis of order 2, and let $F$ be a subset of $A$. The set $A$ is $r$-minimal if $A\setminus F$ is an asymptotic basis if $|F|<r$ and an asymptotic nonbasis if $|F|\geq r$. An asymptotic basis $A$ is $\aleph_{0}$-minimal if $A\setminus F$ is a basis if $F$ is finite and a nonbasis if $F$ is infinite.
Let $A$ be an asymptotic nonbasis of order 2, and let $G$ be a subset of $\mathbf{N}_{0}\setminus A$. The set $A$ is $s$-maximal if $A\cup G$ is an asymptotic nonbasis if $|G|<s$ and an asymptotic basis if $|G|\geq s$. An asymptotic nonbasis $A$ is $\aleph_{0}$-maximal if $A\cup G$ is a nonbasis if $G$ is finite and a basis if $G$ is infinite.
Theorem 8 (Erdős-Nathanson [16]).
Let $Q=\{2q_{k}+1\}_{k=1}^{\infty}$ be a set of odd integers such that $q_{k}\geq 5q_{k-1}+3$ for all $k\geq 2$. Let
$A^{Q}=\bigcup_{k=2}^{\infty}\left(\left[2q_{k-1}+2,q_{k}-q_{k-1}\right]\cup\left[q_{k}+1,q_{k}+q_{k-1}\right]\right).$
  1. (1)
    For all $r\geq 1$, the set $A^{Q}$ is contained in an $r$-minimal basis of order 2.
  2. (2)
    The set $A^{Q}$ is contained in an $\aleph_{0}$-minimal basis of order 2.
  3. (3)
    For all $s\geq 1$, the set $A^{Q}$ is contained in an $s$-maximal nonbasis of order 2.
Theorem 9 (Erdős-Nathanson [16]).
  1. (1)
    There does not exist an asymptotic basis $A=\{a_{i}:i=1,2,3,\ldots\}$ of order 2 such that $A\setminus\{a_{u}:u\in U\}$ is an asymptotic basis of order 2 if $U$ has zero density and an asymptotic nonbasis of order 2 if $U$ has positive density.
  2. (2)
    There does not exist an $\aleph_{0}$-maximal asymptotic nonbasis of order 2.
A set of positive integers is infinitely oscillating if it switches from basis to nonbasis to basis to nonbasis …of order 2 as integers are alternately adjoined to and deleted from the set. Equivalently, let $A$ be a set of positive integers and let $F$ and $G$ be finite sets such that $F\subseteq A$ and $G\subseteq\mathbf{N}\setminus A$. The set $A$ is an infinitely oscillating basis if $(A\setminus F)\cup G$ is an asymptotic basis of order 2 if and only is $|F|\leq|G|$.
Let $B$ be a set of positive integers and let $F$ and $G$ be finite sets such that $G\subseteq B$ and $F\subseteq\mathbf{N}\setminus B$. The set $B$ is an infinitely oscillating nonbasis if $(B\cup F)\setminus G$ is an asymptotic nonbasis of order 2 if and only is $|F|\leq|G|$.
Infinitely oscillating bases exist. The following stronger result is strange and fascinating.
Theorem 10 (Erdős-Nathanson [17]).
There exists a partition of the positive integers $\mathbf{N}$ into two disjoint sets $A$ and $B$ such that $A$ is an infinitely oscillating basis and $B$ is an infinitely oscillating nonbasis.
The proof is nonconstructive.
Problem 9.
Construct a partition of the positive integers $\mathbf{N}$ into two disjoint sets $A$ and $B$ such that $A$ is an infinitely oscillating asymptotic basis of order 2 and $B$ is an infinitely oscillating asymptotic nonbasis of order 2.
Theorem 10 suggests several new problems. If $A$ is an asymptotic basis of order $h$, then $A$ is an asymptotic basis of order $h^{\prime}$ for all $h^{\prime}\geq h$. The exact order of a set $A$ of integers is the least integer $h$ such that $A$ is an asymptotic basis of order $h$.
Problem 10.
Let $h_{0}<h_{1}<h_{2}<\cdots$ be a strictly increasing finite or infinite sequence of nonnegative integers with $h_{0}\geq 2$.
  1. (1)
    Does there exists an asymptotic basis $A$ of order $h_{0}$ and an infinite sequence $(a_{i})_{i=1}^{\infty}$ of elements of $A$ such that the set $A\setminus\{a_{1},a_{2},\ldots,a_{i}\}$ has exact order $h_{i}$ for all $i$?
  2. (2)
    Does there exist an asymptotic basis $A$ of order $h_{0}$ and an infinite sequence $(a_{i})_{i=1}^{\infty}$ of elements of $A$ such that the set $A\setminus\{a_{i}\}$ has exact order $h_{i}$ for all $i$?
6. Sums of finite sets
A central problem in additive number theory is to understand the growth of a sumset $hA$ as $h\rightarrow\infty$.
The $d$-dilation of the set $A$ is $d\ast A=\{da:a\in A\}$. Then $d\ast\mathbf{N}_{0}=\{0,d,2d,3d,\ldots\}$ is the set of all nonnegative multiples of $d$. For infinite sets there is the following result.
Theorem 11 (Nash-Nathanson [45]).
Let $A$ be an infinite set of nonnegative integers with $d_{L}(A)>0$ and $\gcd(A)=d$. There is an integer $h_{0}$ such that $hA\sim d\ast\mathbf{N}_{0}$ for all $h\geq h_{0}$.
For finite sets, a fundamental theorem of additive number theory describes the structure of the $h$-fold sum of a finite set $A$ of integers for all sufficiently large $h$. Every finite set of integers can be “normalized” by a translation and contraction to obtain a set $A$ such that $\min(A)=0$ and $\gcd(A)=1$.
Theorem 12 (Nathanson [46]).
Let $A$ be a normalized finite set of integers with $|A|\geq 2$ and $a^{*}=\max(A)$. Let $|A|=k+1$ and
$h^{*}=(k-1)a^{*}(a^{*}-1).$
There are integers $C$ and $D$ and sets $\mathcal{C}\subseteq[0,C-2]$ and $\mathcal{D}\subseteq[0,D-2]$ such that, if $h\geq h^{*}$, then
$hA=\mathcal{C}\cup[C,ha^{*}-D]\cup(ha^{*}-\mathcal{D}).$
Recently, Granville and Shakan [28] and Granville and Walker [29] obtained the best possible value for $h_{0}$ and the structure of the sets $\mathcal{C}$ and $\mathcal{D}$ in this result.
Theorem 12 extends to linear forms of sums of finite sets of integers. A finite sequence of finite sets of nonnegative integers can be “normalized” by translations and contractions to obtain a sequence $A_{1},\ldots,A_{r}$ such that $0\in A_{i}$ for all $i\in\{1,\ldots,r\}$ and $\gcd\left(\bigcup_{i=1}^{r}A_{i}\setminus\{0\}\right)=1$.
Theorem 13 (Han, Kirfel, and Nathanson [32]).
Let $A_{1},\ldots,A_{r}$ be a normalized sequence of finite sets of integers with $a_{i}^{*}=\max(A_{i})$ for $i\in\{1,\ldots,r\}$. There are integers $C$ and $D$ and finite sets
$\mathcal{C}\subseteq[0,C-2]\operatorname{\qquad\text{and}\qquad}\mathcal{D}\subseteq[0,D-2]$
and there exist integers $h_{1}^{*},\ldots,h_{r}^{*}$ such that, if $h_{i}\geq h_{i}^{*}$ for all $i\in\{1,\ldots,r\}$, then
$h_{1}A_{1}+\cdots+h_{r}A_{r}=\mathcal{C}\cup\left[C,\sum_{i=1}^{r}h_{i}a_{i}^{*}-D\right]\cup\left(\sum_{i=1}^{r}h_{i}a_{i}^{*}-\mathcal{D}\right).$
Problem 11.
Determine the best possible values for the integers $h_{1}^{*},\ldots,h_{r}^{*}$ and the structure of the sets $\mathcal{C}$ and $\mathcal{D}$ in Theorem 13.
Let $A$ be a set of integers. The $h$-fold sumset $hA$ is the set of integers that can be represented as the sum of $h$ elements of $A$, that is, the integers $n$ with $r_{h,A}(n)\geq 1$. For every positive integer $t$, let $(hA)^{(t)}$ be the set of integers that have at least $t$ representations as sums of $h$ elements of $A$, that is, the integers $n$ with $r_{h,A}(n)\geq t$. The structure of the sumsets $(hA)^{(t)}$ has recently been determined.
Theorem 14 (Nathanson [58]).
Let $A$ be a normalized finite set of integers with $a^{*}=\max(A)$. For every positive integer $t$, let
$h_{t}^{*}=(k-1)(ta^{*}-1)a^{*}+1.$
There are nonnegative integers $C_{t}$ and $D_{t}$ and sets $\mathcal{C}_{t}\subseteq[0,C_{t}-2]$ and $\mathcal{D}_{t}\subseteq[0,D_{t}-2]$ such that, if $h\geq h_{t}^{*}$, then
$(hA)^{(t)}=\mathcal{C}_{t}\cup[C_{t},ha^{*}-D_{t}]\cup(ha^{*}-\mathcal{D}_{t}).$
Problem 12.
Determine the best possible values for the integers $h_{t}^{*}$ and the structure of the sets $\mathcal{C}_{t}$ and $\mathcal{D}_{t}$ in Theorem 14.
There is a more subtle additive problem. Color the elements of the set $A$ with $q$ colors, which we call $\{1,\ldots,q\}$. Let $A_{i}$ be the subset of $A$ consisting of all elements $a\in A$ that have color $i$. The sets $A_{1},\ldots,A_{q}$ are pairwise disjoint with $A=\bigcup_{i=1}^{q}A_{i}$. The $q$-tuple $\mathbf{A}=(A_{1},\ldots,A_{q})$ is an ordered partition of $A$. Let $\mathbf{N}_{0}^{q}$ be the set of $q$-tuples of nonnegative integers. For all
$\mathbf{h}=(h_{1},\ldots,h_{q})\in\mathbf{N}_{0}^{q},$
let
$\|\mathbf{h}\|=\sum_{i=1}^{q}h_{i}=h.$
The chromatic subset $\mathbf{h}\cdot\mathbf{A}$ is the set of all integers in the sumset $hA$ that can be represented as the sum of $h$ elements of $A$ with exactly $h_{i}$ elements of color $i$ for all $i\in[1,q]$. Thus, $n\in\mathbf{h}\cdot\mathbf{A}$ if and only if we can write
(2) $n=\sum_{i=1}^{q}\sum_{j_{i}=1}^{h_{i}}a_{i,j_{i}}$
where
(3) $a_{i,j_{i}}\in A_{i}\qquad\text{for all $i\in[1,q]$ and $j_{i}\in[1,h_{i}]$.}$
We refine this problem by allowing elements of $A$ to have more than one color. The subset $A_{i}$ still consists of all elements of $A$ with color $i$, and $A=\bigcup_{i=1}^{q}A_{i}$, but the sets $A_{i}$ are not necessarily pairwise disjoint. The chromatic sumset $\mathbf{h}\cdot\mathbf{A}$ consists of all integers $n$ that have at least one representation of the form (2) and (3). Note that
$h_{i}A_{i}=\left\{\sum_{j_{i}=1}^{h_{i}}a_{i,j_{i}}:a_{i,j_{i}}\in A_{i}\text{ for all }j_{i}\in[1,h_{i}]\right\}$
and so
$\mathbf{h}\cdot\mathbf{A}$ $=(h_{1},\ldots,h_{q})\cdot(A_{1},\ldots,A_{q})$
$=h_{1}A_{1}+\cdots+h_{q}A_{q}.$
This is a homogeneous linear form in the sets $A_{1},\ldots,A_{q}$. For every set $B$ of integers, we also have the inhomogeneous linear form $h_{1}A_{1}+\cdots+h_{q}A_{q}+B$. Han, Kirfel, and Nathanson (Theorem 13) determined the asymptotic structure of homogeneous and inhomogeneous linear forms for all $q$-tuples of finite sets of integers.
The chromatic representation function $r_{\mathbf{A},\mathbf{h}}(n)$ counts the number of colored representations of $n$ of the form (2) and (3), that is, the number of $q$-tuples
$\left(\mathbf{a}_{1},\ldots,\mathbf{a}_{q}\right)\in A_{1}^{h_{1}}\times\cdots\times\cdots A_{q}^{h_{q}}$
where, for all $i\in[1,q]$, the $h_{i}$-tuple $\mathbf{a}_{i}=(a_{i,j_{1}},a_{i,j_{2}},\ldots,a_{i,j_{h_{i}}})\in A_{i}^{h_{i}}$ satisfies
$a_{i,j_{1}}\leq a_{i,j_{2}}\leq\cdots\leq a_{i,j_{h_{i}}}.$
Let
$(\mathbf{h}\cdot\mathbf{A})^{(t)}=\{n\in\mathbf{h}\cdot\mathbf{A}:r_{\mathbf{A},\mathbf{h}}(n)\geq t\}.$
Define a partial order on the set $\mathbf{N}_{0}^{q}$ as follows. For vectors
$\mathbf{h}_{1}=(h_{1,1},\ldots,h_{q,1})\in\mathbf{N}_{0}^{q}\operatorname{\qquad\text{and}\qquad}\mathbf{h}_{2}=(h_{1,2},\ldots,h_{q,2})\in\mathbf{N}_{0}^{q}$
let
$\mathbf{h}_{1}\preceq\mathbf{h}_{2}\qquad\text{ if $h_{1,i}\leq h_{2,i}$ for all $i\in[1,q]$.}$
The following result describes the structure of the chromatic sumset $(\mathbf{h}\cdot\mathbf{A})^{(t)}$ for all positive integers $t$ and all sufficiently large vectors $\mathbf{h}\in\mathbf{N}_{0}^{q}$.
Theorem 15 (Nathanson [57]).
Let $\mathbf{A}=(A_{1},\ldots,A_{q})$ be a normalized $q$-tuple of finite sets of integers. Let $\max(A_{i})=a_{i}^{*}\geq 1$ for all $i\in[1,q]$, and
$\mathbf{a}^{*}=\left(a_{1}^{*},\ldots,a_{q}^{*}\right)\in\mathbf{N}_{0}^{q}.$
For every positive integer $t$, there exist nonnegative integers $c_{t}$ and $d_{t}$, there exist finite sets of nonnegative integers $C_{t}$ and $D_{t}$, and there exists a vector $\mathbf{h}_{t}=(h_{t,1},\ldots,h_{t,q})\in\mathbf{N}_{0}^{q}$ such that, if $\mathbf{h}=(h_{1},\ldots,h_{q})\in\mathbf{N}_{0}^{q}$ and $\mathbf{h}\succeq\mathbf{h}_{t}$, then
$\left(\mathbf{h}\cdot\mathbf{A}\right)^{(t)}=C_{t}\cup\left[c_{t},\mathbf{h}\cdot\mathbf{a}^{*}-d_{t}\right]\cup\left(\mathbf{h}\cdot\mathbf{a}^{*}-D_{t}\right).$
7. Lattice points and abelian semigroups
Let $\mathbf{Z}^{n}$ denote the group of lattice points in $\mathbf{R}^{n}$. The higher-dimensional additive problem is the study of sums of finite sets of lattice points. In discrete geometry, a lattice polytope is the convex hull of a finite set of lattice points. Let $A$ be a finite subset of $\mathbf{Z}^{n}$ that is not contained in a hyperplane of $\mathbf{R}^{n}$, and let $K$ be the convex hull of $A$. For every positive integer $h$, the dilation $h\ast K$ is equal to the $h$-fold sumset $hK$. Ehrhart  [12, 13, 14] proved that there is a polynomial $e_{K}(z)$ of degree $n$ such the number of lattice points in the sumset $hK$, that is, $\operatorname{\text{card}}(hK\cap\mathbf{Z}^{n})$, is exactly $e_{K}(h)$ for all sufficiently large $h$. We have $A\subseteq K$ and so $hA\subseteq hK\cap\mathbf{Z}^{n}$. It follows that
$\operatorname{\text{card}}(hA)\leq\operatorname{\text{card}}(hK\cap\mathbf{Z}^{n})=e_{K}(h).$
Khovanskii [39, 40] proved that there is a polynomial $f_{A}(z)$ of degree $n$ such that
$\operatorname{\text{card}}(hA\cap\mathbf{Z}^{n})=f_{A}(h)$
for all sufficiently large $h$ and that the sumset $hA$ contains every lattice point of $hK$ that is not “close” to the boundary of the polytope $hK$.
Problem 13.
Describe the structure of the set of lattice points in the “boundary layer” of points in $(hK\cap\mathbf{Z}^{n})\setminus hA$.
Now let $X$ be any additive abelian semigroup. Khovanskii proved that if $A$ and $B$ are finite subsets of $X$, then there is a polynomial $g_{A,B}(z)$ such that
$\operatorname{\text{card}}(hA+B)=g_{A,B}(h)$
for all sufficiently large $h$.
Nathanson extended this result to linear forms of finite sequences of finite subsets of $X$.
Theorem 16 (Nathanson [52]).
Let $A_{1},\ldots,A_{\ell}$, and $B$ be finite subsets of an additive abelian semigroup $X$. There exists a polynomial $p_{A_{i},B}(z_{1},\ldots,z_{\ell})$ and an integer $h^{*}=h^{*}(A_{1},\ldots,A_{\ell},B)$ such that, if $h_{1},\ldots,h_{\ell}$ are integers with $\min(h_{i})\geq h^{*}$, then
$\operatorname{\text{card}}\left(h_{1}A_{1}+\cdots+h_{\ell}A_{\ell}+B)\cap\mathbf{Z}^{n}\right)=p_{A_{i},B}(h_{1},\ldots,h_{\ell}).$
The proof uses the Hilbert polynomial for finitely generated graded rings. Nathanson and Ruzsa [66] gave an elementary proof of the special case $B=\{0\}$.
8. Approximate groups
Let $A$ be a nonempty subset of an additive abelian group $G$, and let $r$ and $\ell$ be positive integers. The set $A$ is an $(r,\ell)$-approximate group if there exists a set $X\subseteq G$ such that
(4) $|X|\leq\ell$
and
(5) $rA\subseteq X+A.$
The idea of an approximate group evolved from Freiman’s inverse theorem in additive number theory (Freiman [24, 25, 26, 27], Nathanson [50]). This definition of approximate group is less restrictive than the original definition, which appears in Tao [76] and which has been extensively investigated (for example, by Breuillard, Green, and Tao [2, 3, 4, 30, 31] and Pyber and Szabo [69]). Helfgott [33] is a recent survey of related problems on growth in groups.
The set $A$ is an asymptotic $(r,\ell)$-approximate group if, for all sufficiently large $h$, the $h$-fold sumset of $A$ is an $(r,\ell)$-approximate group. This means that there exists an integer $h_{0}(A)$ such that, for each integer $h\geq h_{0}(A)$, there is a set $X_{h}\subseteq G$ such that
(6) $|X_{h}|\leq\ell$
and
(7) $rhA\subseteq X_{h}+hA.$
Theorem 17 (Nathanson [55]).
Every nonempty finite subset of an abelian group is an asymptotic approximate group.
Problem 14.
Theorem 17 is a very general result. It would be of interest to obtain explicit estimates for $r$, $h$, and $\ell$ for finite sets of lattice points.
9. Sidon sets and $B_{h}$-sets
The Erdős-Turán conjecture asserts the unboundedness of the representation function of an asymptotic basis. In the opposite direction, we consider sets with sumsets whose representation functions are bounded. The extreme case is a Sidon set, which is a finite or infinite subset $A$ of the nonnegative integers (or of any additive abelian semigroup) such that $r_{2,A}(n)\leq 1$ for all $n\in\mathbf{N}_{0}$. More generally, a $B_{h}$-set is a finite or infinite set $A$ of the nonnegative integers such that $r_{h,A}(n)\leq 1$ for all $n\in\mathbf{N}_{0}$. A Sidon set is a $B_{2}$-set. A $B_{h}[g]$-set is a finite or infinite set $A$ of nonnegative integers such that $r_{h,A}(n)\leq g$ for all $n\in\mathbf{N}_{0}$. There is a vast literature on these sets (O’Bryant [67]) and many open problems. For example, what is the cardinality of the largest Sidon set contained in the interval $[1,n]$? The current best result is due to Balogh, Füredi, and Roy [1].
Of recent interest are perturbations of Sidon sets and $B_{h}$-sets. Let $\mathbf{F}$ be a field with a nontrivial absolute value |||\ |. Let $I$ be a nonempty set, and let $\{a_{i}:i\in I\}$ and $\{b_{i}:i\in I\}$ be sets of elements of the field $\mathbf{F}$. Let $\{\varepsilon_{i}:i\in I\}$ be a set of positive real numbers. The set $\{b_{i}:i\in I\}$ is an $\varepsilon$-perturbation of the set $\{a_{i}:i\in I\}$ if $|b_{i}-a_{i}|<\varepsilon_{i}$ for all $i\in I$.
Theorem 18 (Nathanson [59]).
Let $\mathbf{F}$ be a field of characteristic 0 with a nontrivial absolute value. Let $\{\varepsilon_{i}:i\in\mathbf{N}\}$ be a set of positive real numbers. For every subset $\{a_{i}:i\in\mathbf{N}\}$ of $\mathbf{F}$, there is a $B_{h}$-set $\{b_{i}:i\in\mathbf{N}\}$ in $\mathbf{F}$ that is an $\epsilon$-perturbation of $A$.
Problem 15.
Let $A$ be an uncountably infinite set. Is there an $\varepsilon$-perturbation of $A$ that is an $h$-Sidon set?
The greedy $B_{h}$-set is the $B_{h}$-set constructed inductively by letting $a_{0}(h)=0$, $a_{1}(h)=1$, and, given the $B_{h}$-set $\{a_{0}(h),a_{1}(h),\ldots,a_{k}(h)\}$, choosing $a_{k+1}(h)>a_{k}(h)$ as the smallest integer such that $\{a_{0}(h),a_{1}(h),\ldots,a_{k}(h),a_{k+1}(h)\}$ is a $B_{h}$-set. We know very little about the structure of these much investigated, but still mysterious, sets constructed with the greedy algorithm. We have $a_{0}(h)=0$, $a_{1}(h)=1$, and $a_{2}(h)=h+1$. Recently, Nathanson [64] proved that
$a_{3}(h)=h^{2}+h+1$
for all $h\geq 1$. Nathanson and O’Bryant [65] proved that
$a_{4}(h)=\begin{cases}\left(h+1\right)^{3}/2&\text{if $h$ is odd}\\ \left(h^{3}+2h^{2}+3h+2\right)/2&\text{if $h$ is even.}\end{cases}$
O’Bryant [obry24] has determined upper and lower bounds for $a_{5}(h)$.
Problem 16.
Compute $a_{5}(h)$ and determine if $a_{k}(h)$ is a quasi-polynomial for all $h\geq 5$.
For other recent work on Sidon sets, see Nathanson [59, 60, 62].
10. Sidon sets for linear forms
Let $\mathbf{F}$ be a field and let $h$ be a positive integer. We consider linear forms
(8) $\varphi(x_{1},\ldots,x_{h})=c_{1}x_{1}+\cdots+c_{h}x_{h}$
where $c_{i}\in\mathbf{F}$ for all $i\in\{1,\ldots,h\}$. Let $A$ be a nonempty subset of a vector space $V$ over the field $\mathbf{F}$. The $\varphi$-image of $A$ is the sum of dilates
$\varphi(A)$ $=\left\{\varphi(a_{1},\ldots,a_{h}):(a_{1},\ldots,a_{h})\in A^{h}\right\}$
$=\left\{c_{1}a_{1}+\cdots+c_{h}a_{h}:(a_{1},\ldots,a_{h})\in A^{h}\right\}$
$=c_{1}\ast A+\cdots+c_{h}\ast A.$
A nonempty subset $A$ of $V$ is a Sidon set for the linear form $\varphi$ or, simply, a $\varphi$-Sidon set if it satisfies the following property: For all $h$-tuples $(a_{1},\ldots,a_{h})\in A^{h}$ and $(a^{\prime}_{1},\ldots,a^{\prime}_{h})\in A^{h}$, if
$\varphi(a_{1},\ldots,a_{h})=\varphi(a^{\prime}_{1},\ldots,a^{\prime}_{h})$
then $(a_{1},\ldots,a_{h})=(a^{\prime}_{1},\ldots,a^{\prime}_{h})$, that is, $a_{i}=a^{\prime}_{i}$ for all $i\in\{1,\ldots,h\}$.
There is a simple obstruction to the existence of $\varphi$-Sidon sets with more than one element. For every nonempty subset $I$ of $\{1,\ldots,h\}$, define the subset sum
(9) $s(I)=\sum_{i\in I}c_{i}.$
Let $s({\emptyset})=0$. Suppose there exist disjoint subsets $I_{1}$ and $I_{2}$ of $\{1,\ldots,h\}$ with $I_{1}$ and $I_{2}$ not both empty such that
(10) $s({I_{1}})=\sum_{i\in I_{1}}c_{i}=\sum_{i\in I_{2}}c_{i}=s({I_{2}}).$
Let $I_{3}=\{1,\ldots,h\}\setminus(I_{1}\cup I_{2})$. Let $A$ be a subset of $V$ with $|A|\geq 2$. Choose vectors $u,v,w\in A$ with $u\neq v$, and define
[…]
We have
$(a_{1},\ldots,a_{h})\neq(a^{\prime}_{1},\ldots,a^{\prime}_{h})$
because $I_{1}\cup I_{2}\neq\emptyset$ and $a_{i}\neq a^{\prime}_{i}$ for all $i\in I_{1}\cup I_{2}$.
The sets $I_{1}$, $I_{2}$, $I_{3}$ are pairwise disjoint. Condition (10) implies
$\varphi(a_{1},\ldots,a_{h})$ $=\left(\sum_{i\in I_{1}}c_{i}\right)u+\left(\sum_{i\in I_{2}}c_{i}\right)v+\left(\sum_{i\in I_{3}}c_{i}\right)w$
$=\left(\sum_{i\in I_{2}}c_{i}\right)u+\left(\sum_{i\in I_{1}}c_{i}\right)v+\left(\sum_{i\in I_{3}}c_{i}\right)w$
$=\varphi(a^{\prime}_{1},\ldots,a^{\prime}_{h})$
and so $A$ is not a $\varphi$-Sidon set. We say that the linear form (8) has property $N$ if there do not exist disjoint subsets $I_{1}$ and $I_{2}$ of $\{1,\ldots,h\}$ that satisfy condition (10) with $I_{1}$ and $I_{2}$ not both empty.
Sidon sets for linear forms were introduced by Nathanson, who proved the following theorems.
Theorem 19 (Nathanson [61]).
Let $\mathbf{F}$ be a field, let $V$ be an infinite vector space over the field $\mathbf{F}$, and let $X$ be an infinite subset of $V$. Let $\varphi(x_{1},\ldots,x_{h})=\sum_{i=1}^{h}c_{i}x_{i}$ be a linear form with nonzero coefficients $c_{i}\in\mathbf{F}$. The following are equivalent:
  1. (i)
    The set $X$ contains an infinite $\varphi$-Sidon set $A$.
  2. (ii)
    The set $X$ contains a $\varphi$-Sidon set $A$ with $|A|\geq 2$.
  3. (iii)
    The linear form $\varphi$ has property $N$.
Theorem 20 (Nathanson [61]).
Let $\mathbf{F}$ be a field with a nontrivial absolute value and let $V$ be a vector space over $\mathbf{F}$ that has a norm with respect to the absolute value on $\mathbf{F}$. Let $\varepsilon=\{\varepsilon_{k}:k=1,2,3,\ldots\}$ be a set of positive real numbers. Let $\varphi$ be a linear form with coefficients in $\mathbf{F}$ that has property $N$. For every set $B=\{b_{k}:k=1,2,3,\ldots\}$ of vectors in $V$, there is a $\varphi$-Sidon set $A=\{a_{k}:k=1,2,3,\ldots\}$ of vectors in $V$ such that
(11) $\|a_{k}-b_{k}\|<\varepsilon_{k}$
for all $k=1,2,3,\ldots$.
Theorem 21 (Nathanson [61]).
Let $\mathbf{F}$ be a field with a nontrivial absolute value, and let $\varphi$ be a linear form with coefficients in $\mathbf{F}$ that has property $N$. Let $V$ be a vector space over $\mathbf{F}$ that has a norm with respect to absolute value on $\mathbf{F}$. For every set $B=\{b_{k}:k=1,2,3,\ldots\}$ of vectors in $V$, there exists a $\varphi$-Sidon set $A=\{a_{k}:k=1,2,3,\ldots\}$ in $V$ such that
$\lim_{k\rightarrow\infty}\|a_{k}-b_{k}\|=0.$
Let $\mathbf{P}=\{2,3,5,\ldots\}$ be the set of prime numbers. For every prime number $p$, let $|\ |_{p}$ be the usual $p$-adic absolute value with $|p|_{p}=1/p$.
Theorem 22 (Nathanson [61]).
Let $\varphi$ be a linear form with rational coefficients that satisfies property $N$. Let $\{\varepsilon_{k}:k=1,2,3,\ldots\}$ be a sequence of positive real numbers and let $\{p_{k}:k=1,2,3,\ldots\}$ be a sequence of prime numbers. For every sequence of integers $B=\{b_{k}:k=1,2,3,\ldots\}$, there exists a strictly increasing sequence of positive integers $A=\{a_{k}:k=1,2,3,\ldots\}$ such that $A$ is a $\varphi$-Sidon set and
$|a_{k}-b_{k}|_{p_{j}}<\varepsilon_{k}$
for all $k\in\mathbf{N}$ and $j\in\{1,\ldots,k\}$.
Problem 17.
Let $\varphi=\sum_{i=1}^{h}c_{i}x_{i}$ be a linear form with integer coefficients. Let $\mathbf{P}$ be the set of prime numbers and let $A=\{\log p:p\in\mathbf{P}\}$. Consider the $h$-tuple $(p_{1},\ldots,p_{h})\in\mathbf{P}^{h}$ of not necessarily distinct prime numbers, and let $\mathbf{P}_{0}=\{p\in\mathbf{P}:p=p_{i}\text{ for some }i\in\{1,\ldots,h\}\}$. For each $p\in\mathbf{P}_{0}$, let
$I_{p}=\{i\in\{1,\ldots,h\}:p_{i}=p\}\operatorname{\qquad\text{and}\qquad}s(I_{p})=\sum_{i\in I_{p}}c_{i}.$
We have
$\varphi(\log p_{1},\ldots,\log p_{h})=\sum_{i=1}^{h}c_{i}\log p_{i}=\sum_{p\in\mathbf{P}_{0}}s(I_{p})\log p=\log\prod_{p\in\mathbf{P}_{0}}p^{S(I_{p})}.$
If the linear form $\varphi$ satisfies property $N$, then, by the fundamental theorem of arithmetic, the set $A=\{\log p:p\in\mathbf{P}\}$ is a $\varphi$-Sidon set.
For the linear form $\psi=x_{1}+\cdots+x_{h}$, Ruzsa [71] used the set $A$ to construct large classical Sidon sets of positive integers . Are such constructions also possible for $\varphi$-Sidon sets of positive integers?
Problem 18.
Let $A=\{a_{k}:k=1,2,3,\ldots\}$ and $B=\{b_{k}:k=1,2,3,\ldots\}$ be sequences of integers. The set $A$ is a polynomial perturbation of $B$ if
$|a_{k}-b_{k}|<k^{r}$
for some $r>0$ and all $k\geq k_{0}$. The set $A$ is a bounded perturbation of $B$ if
$|a_{k}-b_{k}|<m_{0}$
for some $r>0$ and all $k\geq k_{0}$.
Let $\varphi$ be a linear form with integer coefficients that satisfies condition $N$. Let $B$ be a set of integers. Does there exist a $\varphi$-Sidon set of integers that is a polynomial perturbation of $B$?
Does there exist a $\varphi$-Sidon set of integers that is a bounded perturbation of $B$?
Problem 19.
Let $\varphi$ be a linear form with integer coefficients that satisfies condition $N$. For every positive integer $n$, determine the cardinality of the largest $\varphi$-Sidon subset of $\{1,2,\ldots,n\}$.
Problem 20.
There exists $c>0$ such that, for every positive integer $n$, there is a classical Sidon set $A\subseteq\{1,\ldots,n\}$ with $A(n)\geq c\sqrt{n}$. However, there is no infinite classical Sidon set $A$ of positive integers such that $A(n)\geq c\sqrt{n}$ for some $c>0$ and all $n\geq n_{0}$. Indeed, Erdős (in Stöhr [72]) proved that every infinite classical Sidon set satisfies
$\liminf_{n\rightarrow\infty}A(n)\sqrt{\frac{\log n}{n}}\ll 1.$
Are there analogous lower bounds for infinite $\varphi$-Sidon sets of positive integers associated with binary linear forms $\varphi=c_{1}x_{1}+c_{2}x_{2}$ or with linear forms $\varphi=\sum_{i=1}^{h}c_{i}x_{i}$ for $h\geq 3$?
Problem 21.
Consider sets of integers. One might expect that the elements of a set $A$ of integers that is “sufficiently random” or “in general position” will be a classical Sidon set, that is, will not contain a nontrivial solution of the equation $x_{1}+x_{2}=x_{3}+x_{4}$. Equivalently, the set $A$ will be one-to-one (up to transposition) on the function $f(x_{1},x_{2})=x_{1}+x_{2}$. There is nothing special about the function $x_{1}+x_{2}$. One could ask if $A$ is one-to-one (up to permutation) on some symmetric function, or one-to-one on a function that is not symmetric. The functions considered in this paper are linear forms in $h$ variables.
Conversely, given the set $A$ of integers, we can ask what are the functions (in some particular set $\mathcal{F}$ of functions) with respect to which the set $A$ is one-to-one. This inverse problem is considered in Nathanson [61].